# Exploring Partial Multi-Label Learning via Integrating Semantic Co-occurrence Knowledge

### v1

@@ III-B Cross-Modality Fusion Module
The cross-modality fusion module employs a fusion strategy to optimize label confidence, that is, the selected visual features and the optimized label confidence are used to train the classifier. This module is specifically designed for deep, interactive fusion of information from multiple modalities, such as textual and visual data. It considers the local similarities between samples and global correlations among labels. This fusion strategy enhances the model’s performance when dealing with partial multi-label data, especially in the presence of noisy labels. In this way, the algorithm can estimate the confidence of each label more accurately, thereby providing the model with a more reliable learning basis.
In the problem of partial multi-label learning, each instance may be associated with a set of candidate labels, among which some may be irrelevant. We construct an objective function by considering global instance similarity and local label relevance to compute the confidence of each label.
@@
Calculating label confidence is essential for evaluating the dependability of different labels when resolving such problems. Moreover, evaluating the global relevance among labels is also crucial since it can reveal potential interconnections among them. We employ the Pearson correlation coefficient to calculate the label correlation, denoted as $r$, to quantify the inter-label relevance. This way, we are able to effectively identify and understand the interrelationships among labels, thereby achieving more accurate predictions in multi-label classification tasks. However, it is essential to recognize that this coefficient does not convey information regarding the causal relationships between labels. Even if there is a high correlation between two labels, it does not directly imply that one label causes changes in the other.
@@ III-C Intrinsic Semantic Augmentation Strategy
The proposed intrinsic semantic augmentation strategy aims to boost model performance in situations where labeling is incomplete. By deeply understanding and leveraging semantic information, particularly the co-occurrence relationships between labels, Intrinsic semantic augmentation strategy enables the model to more effectively capture the inherent connections between labels. This approach acknowledges that models can achieve better results in MLR tasks by fully exploiting these underlying semantic associations, even when faced with partial or missing labels.
@@
In multi-label classification, each instance may belong to multiple categories. Existing loss functions typically reweight each label independently, without considering the inter-label and inter-instance correspondences. We propose a triple transformation strategy to enhance the model’s understanding of images to address this issue. By subjecting the input image to weak, medium (i.e., the untransformed original image), and strong transformations, we strive for the model to learn richer information from disturbances of varying degrees. During each augmentation, weak and strong transformations are applied by randomly selecting one out of one and five types of transformations, respectively. These three transformations correspond to different semantic distributions: $P\left(Y|\omega\left(X^{-}\right)\right)$, $P\left(Y|\theta\left(X\right)\right)$, and $P\left(Y|\Omega\left(X^{+}\right)\right)$, all in $\mathbb{R}^{n\times 1}$. The weak transformation $\omega\left(X^{-}\right)$ is designed to make subtle adjustments to the image, such as random cropping, horizontal flipping, and color jittering, preserving most of the original semantic information, thereby aiding the model in identifying and learning the core elements within the image. The medium transformation $\theta\left(X\right)$ provides a robust baseline, ensuring that the model can accurately capture the fundamental features of the image without being disrupted by the transformation. On the other hand, the strong transformation $\Omega\left(X^{+}\right)$ modifies the image in a more aggressive manner, such as random rotation, mixup, and cutmix, introducing greater sample diversity and promoting the model’s ability to extract features in a more flexible and robust way. Subsequently, we construct a confident label set $\mathcal{C}\left(x\right)$ to enhance consistency among the three transformations. For each label category, a dynamic thresholding strategy is implemented: A label $c$ is only incorporated into the consistency loss computation when its probability surpasses the predefined threshold $\mathcal{K}$. This mechanism effectively mitigates interference from uncertain labels while ensuring consistent optimization of identical labels across multiple transformations. It is worth noting that although the weak transformation $\omega\left(X^{-}\right)$ and the strong transformation $\Omega\left(X^{+}\right))$ perturb the images to different extents, both are processed by the same visual encoder. This ensures that the features are mapped to a unified space, thereby facilitating contrastive learning and consistency constraints. Sharing the encoder also promotes parameter sharing and regularization, enabling the model to focus on capturing the core semantics of the images and thereby enhancing its robustness and generalization capabilities.
@@
In practice, data transformations may not yield consistent outcomes. This implies that for learning algorithms, the introduction of adaptability strategies and semantic structure enhancement can help the model better understand and deal with inconsistencies in data structures. Considering the adoption of three unique transformation mechanisms, we perform internal knowledge transfer through a self-distillation strategy to enhance the model’s efficacy, leveraging knowledge accumulated at different stages of the training process. By optimizing the Kullback-Leibler divergence, we guide the semantic distribution calibration of transformed images to improve their distribution:
@@
These three transformations complement each other to form a comprehensive learning strategy aimed at enhancing the model’s overall understanding and adaptability to data. Subtle adjustments in the weak transformations enable the model to identify and learn the core elements of the image. The medium transformations provide a reliable baseline, ensuring the model can accurately capture the essential features of the image without any interference from transformations. On the other hand, strong transformations introduce sample diversity through larger image modifications, encouraging more robust and flexible feature extraction capabilities.
Pareto optimality emphasizes the trade-offs and balances in multi-objective optimization to avoid weakening other objectives while improving one [42, 43]. Inspired by this, we combine three different losses and utilize Pareto optimization theory to balance multiple loss functions. We place the multiple losses into a list, calculate the Pareto frontier, and update the loss weights.
@@ IV-A4 Evaluation
[…]
${\rm mAP}=\frac{1}{\left|\mathcal{Y}\right|}\underset{i=1}{\overset{\left|\mathcal{Y}\right|}{\mathop{\sum}}}\,\text{A}{{\text{P}}_{i}},\$ (11)
@@
[…]
@@ IV-C Ablation Study
The proposed SCINet exhibits substantial enhancements in dataset performance across various datasets. SCINet attains a 7.88% enhancement in mAP relative to the baseline model on the VOC2012 dataset. In the COCO2014 dataset with singular labels, the mAP enhancement is 4.86%, whereas in the CUB dataset, the mAP improvement attains 7.31%, underscoring SCINet’s distinct superiority in fine-grained classification tasks. Furthermore, on the VOC2007 dataset with incomplete labels, SCINet’s avg. mAP rises by 9.02%, whereas on the COCO2014 dataset with incomplete labels, the mAP increase is 4.20%. The results demonstrate that SCINet can substantially improve model classification performance across diverse label contexts. SCINet significantly outperforms the baseline model across many datasets due to the thorough integration of components including the bi-dominant prompter module, cross-modality fusion module, and intrinsic semantic augmentation strategy. SCINet’s avg. mAP across all datasets exhibits a 6.66% enhancement relative to the baseline model, underscoring its robust competitiveness and extensive adaptability in multi-label classification tasks. This further substantiates SCINet’s capacity to sustain outstanding performance throughout varied and intricate tasks, as well as its considerable advantage in numerous labeling contexts. Furthermore, this study examined the layer depth findings of graph convolutional networks and studied the effects of single label configurations and label correlation methods on performance, as seen in Fig. 4 and Table IV.
@@ IV-D Adaptive Dynamic Threshold Analysis
We evaluated the model’s adaptive dynamic threshold, as shown in Fig. 10. By adjusting the threshold parameter $\mathcal{K}$, we observed its impact on the model’s attention distribution and performance. For larger and more complex datasets, such as COCO2014, moderately lowering the temperature value helped the model achieve better performance on key metrics. Notably, when $\mathcal{K}=0.3$, the model demonstrated the best balance between precision and recall, suggesting that lower threshold values can effectively enhance the model’s ability to recognize minority classes, which is particularly crucial when dealing with complex datasets. This indicates that dynamically adjusting the threshold $\mathcal{K}$ plays a significant role in optimizing performance on large-scale, complex datasets.
@@ IV-E Visualization of the Quantity of Learnable Prompts
Fig. 8 illustrates the effect of the number of learnable prompts on object recognition performance across various scenes. As the prompt length increases, the model’s false detection rate (highlighted in red) also increases. For example, in Fig. 8(a), when the prompt length is 4 (first row) and 8 (second row), the detection is entirely accurate. However, when the prompt length increases to 32 (fourth row), false detections become significantly more frequent, with multiple incorrect annotations such as “Bicycle: 0.5502” and “Bottle: 0.8534.” This suggests that, particularly in complex scenes, while increasing prompt length can enhance the model’s detection capabilities, it may also introduce more errors. A similar trend is observed in Figs. 8(b) and 8(d).In Fig. 8(c), when the prompt length is 4 (first row), the model successfully detects the correct object label (“Banana: 0.5610”) but also produces false labels, such as “Dining Table: 0.5085.” This indicates that, with shorter prompts, the model can correctly identify some objects while still suffering from false detections. However, when the prompt length increases to 8, 16, and 32 (second to fourth rows), the detection results exhibit notable errors, including high-confidence false labels like “Chair: 0.8581,” “TV monitor: 0.9042,” and “Dining Table: 0.7844.” These results suggest that in complex scenes containing numerous small objects, shorter prompt lengths help maintain the model’s object recognition capabilities, whereas longer prompt lengths may incorporate additional features that lead to a significant increase in the false detection rate.
@@ IV-G Limitation and Future Work
While SCINet demonstrates state-of-the-art performance in partial multi-label learning tasks, it still presents certain limitations. Visual analysis reveals that increasing the length of learnable prompts can enhance detection capabilities; however, this often comes at the cost of a higher false detection rate, particularly in complex scenes with numerous small objects. This suggests a trade-off. Longer prompts may introduce additional, potentially noisy features, thereby undermining the model’s robustness and generalization ability, especially in scenarios characterized by occlusions or background clutter.
@@ V Conclusions
In this study, we introduce SCINet, a novel framework designed to tackle the challenges of partial multi-label learning tasks. At the core of SCINet is the use of structured semantic prior from multi-modal representation learning to make up for the lack of supervisory information by introducing a semantic co-occurrence insight network. SCINet combines text and visual features through a cross-modal fusion module, enhancing the model’s contextual understanding, feature extraction efficiency, and classification accuracy. Moreover, through the intrinsic semantic augmentation strategy, SCINet not only learns the features of the data but also captures its semantic meanings. This enables the model to excel in handling rare or ambiguous categories and improves its ability to generalize and manage the complexity and diversity of the real world.

### v2

@@ III-B Cross-Modality Fusion Module
The cross-modality fusion module employs a fusion strategy to optimize label confidence, utilizing both the selected visual features and the optimized label confidence to train the classifier. This module is specifically designed for deep, interactive fusion of information from multiple modalities, such as textual and visual data. It considers the local similarities between samples and global correlations among labels. This fusion strategy enhances the model’s performance when handling partial multi-label data, particularly in the presence of noisy labels. In this way, the algorithm can estimate the confidence of each label more accurately, thereby providing the model with a more reliable learning basis.
In the problem of PML, each instance may be associated with a set of candidate labels, some of which may be irrelevant to the problem. We construct an objective function that considers both global instance similarity and local label relevance to compute the confidence of each label.
@@
Calculating label confidence is crucial for assessing the reliability of different labels when resolving such issues. Moreover, evaluating the global relevance of labels is also crucial, as it can reveal potential interconnections among them. We employ the Pearson correlation coefficient to calculate the label correlation, denoted as $r$, to quantify the inter-label relevance. This approach enables us to effectively identify and understand the interrelationships among labels, thereby achieving more accurate predictions in multi-label classification tasks. However, it is essential to recognize that this coefficient does not convey information regarding the causal relationships between labels. Even if there is a high correlation between two labels, it does not directly imply that one label causes changes in the other.
@@ III-C Intrinsic Semantic Augmentation Strategy
The proposed intrinsic semantic augmentation strategy aims to boost model performance in situations where labeling is incomplete. By deeply understanding and leveraging semantic information, particularly the co-occurrence relationships between labels, the Intrinsic semantic augmentation strategy enables the model to capture the inherent connections between labels more effectively. This approach acknowledges that models can achieve better results in MLR tasks by fully exploiting these underlying semantic associations, even when faced with partial or missing labels.
@@
In multi-label classification, each instance may belong to multiple categories. Existing loss functions typically reweight each label independently, without considering the inter-label and inter-instance correspondences. We propose a triple transformation strategy to enhance the model’s understanding of images to address this issue. By subjecting the input image to weak, medium (i.e., the untransformed original image), and strong transformations, we strive for the model to learn richer information from disturbances of varying degrees. During each augmentation, weak and strong transformations are applied by randomly selecting one out of one and five types of transformations, respectively. These three transformations correspond to different semantic distributions: $P\left(Y|\omega\left(X^{-}\right)\right)$, $P\left(Y|\theta\left(X\right)\right)$, and $P\left(Y|\Omega\left(X^{+}\right)\right)$, all in $\mathbb{R}^{n\times 1}$. The weak transformation $\omega\left(X^{-}\right)$ is designed to make subtle adjustments to the image, such as random cropping, horizontal flipping, and color jittering, preserving most of the original semantic information, thereby aiding the model in identifying and learning the core elements within the image. The medium transformation $\theta\left(X\right)$ provides a robust baseline, ensuring that the model can accurately capture the fundamental features of the image without being disrupted by the transformation. On the other hand, the strong transformation $\Omega\left(X^{+}\right)$ modifies the image in a more aggressive manner, such as random rotation, mixup, and cutmix, introducing greater sample diversity and promoting the model’s ability to extract features in a more flexible and robust way. Subsequently, we construct a confident label set $\mathcal{C}\left(x\right)$ to enhance consistency among the three transformations. For each label category, a dynamic thresholding strategy is implemented: A label $c$ is only incorporated into the consistency loss computation when its probability surpasses the predefined threshold $\mathcal{K}$. This mechanism effectively mitigates interference from uncertain labels while ensuring consistent optimization of identical labels across multiple transformations. It is worth noting that although the weak transformation $\omega\left(X^{-}\right)$ and the strong transformation $\Omega\left(X^{+}\right)$ perturb the images to different extents, both are processed by the same visual encoder. This ensures that the features are mapped to a unified space, thereby facilitating contrastive learning and consistency constraints. Sharing the encoder also promotes parameter sharing and regularization, enabling the model to focus on capturing the core semantics of the images and thereby enhancing its robustness and generalization capabilities.
@@
In practice, data transformations may not yield consistent outcomes. This implies that for learning algorithms, the introduction of adaptability strategies and semantic structure enhancements can help the model better understand and deal with inconsistencies in data structures. Considering the adoption of three unique transformation mechanisms, we perform internal knowledge transfer through a self-distillation strategy to enhance the model’s efficacy, leveraging knowledge accumulated at different stages of the training process. By optimizing the Kullback-Leibler divergence, we guide the semantic distribution calibration of transformed images to improve their distribution:
@@
These three transformations complement each other to form a comprehensive learning strategy that enhances the model’s overall understanding and adaptability to data. Subtle adjustments in the weak transformations enable the model to identify and learn the core elements of the image. The medium transformations provide a reliable baseline, ensuring the model can accurately capture the essential features of the image without any interference from transformations. On the other hand, vigorous transformations introduce sample diversity through larger image modifications, encouraging more robust and flexible feature extraction capabilities.
Pareto optimality emphasizes the trade-offs and balances in multi-objective optimization to avoid weakening other objectives while improving one [13, 43]. Inspired by this, we combine three different losses and utilize Pareto optimization theory to balance multiple loss functions. We compile the multiple losses into a list, calculate the Pareto frontier, and update the loss weights accordingly.
@@ IV-A4 Evaluation
${\rm A{{P}_{i}}}=\frac{1}{\left|{{G}_{i}}\right|}\underset{k=1}{\overset{n}{\mathop{\sum}}}\,{{P}_{k}\times re{{l}_{k}}},\\$ (10)
${\rm mAP}=\frac{1}{\left|\mathcal{Y}\right|}\underset{i=1}{\overset{\left|\mathcal{Y}\right|}{\mathop{\sum}}}\,\text{A}{{\text{P}}_{i}},\\$ (11)
@@
${\rm OF1}=\frac{2\times TP}{2\times TP+FP+FN},\\$ (12)
${\rm CF1}=\frac{1}{K}\underset{i=1}{\overset{K}{\mathop{\sum}}}\,\frac{2\times T{{P}_{i}}}{2\times T{{P}_{i}}+F{{P}_{i}}+F{{N}_{i}}},\\$ (13)
@@ IV-C Ablation Study
The proposed SCINet exhibits substantial enhancements in dataset performance across various datasets. SCINet attains a 7.88% enhancement in mAP relative to the baseline model on the VOC2012 dataset. In the COCO2014 dataset with singular labels, the mAP enhancement is 4.86%, whereas in the CUB dataset, the mAP improvement attains 7.31%, underscoring SCINet’s distinct superiority in fine-grained classification tasks. Furthermore, on the VOC2007 dataset with incomplete labels, SCINet’s avg. mAP rises by 9.02%, whereas on the COCO2014 dataset with incomplete labels, the mAP increase is 4.20%. The results demonstrate that SCINet can substantially improve model classification performance across diverse label contexts. SCINet significantly outperforms the baseline model across many datasets due to the thorough integration of components, including the bi-dominant prompter module, cross-modality fusion module, and intrinsic semantic augmentation strategy. SCINet’s avg. mAP across all datasets exhibits a 6.66% enhancement relative to the baseline model, underscoring its robust competitiveness and extensive adaptability in multi-label classification tasks. This further substantiates SCINet’s capacity to sustain outstanding performance throughout varied and intricate tasks, as well as its considerable advantage in numerous labeling contexts. Furthermore, this study examined the layer depth findings of graph convolutional networks and studied the effects of single-label configurations and label correlation methods on performance, as seen in Fig. 4 and Table IV.
@@ IV-D Adaptive Dynamic Threshold Analysis
We evaluated the model’s adaptive dynamic threshold, as shown in Fig. 10. By adjusting the threshold parameter $\mathcal{K}$, we observed its impact on the model’s attention distribution and performance for larger and more complex datasets, such as COCO 2014. Moderately lowering the temperature value improved the model’s performance on key metrics. Notably, when $\mathcal{K}=0.3$, the model demonstrated the best balance between precision and recall, suggesting that lower threshold values can effectively enhance the model’s ability to recognize minority classes, which is particularly crucial when dealing with complex datasets. This indicates that dynamically adjusting the threshold $\mathcal{K}$ plays a significant role in optimizing performance on large-scale, complex datasets.
@@ IV-E Visualization of the Quantity of Learnable Prompts
Fig. 8 illustrates the effect of the number of learnable prompts on object recognition performance across various scenes. As the prompt length increases, the model’s false detection rate (highlighted in red) also increases. For example, in Fig. 8(a), when the prompt length is 4 (first row) and 8 (second row), the detection is entirely accurate. However, when the prompt length increases to 32 (fourth row), false detections become significantly more frequent, with multiple incorrect annotations such as “Bicycle: 0.5502” and “Bottle: 0.8534.” This suggests that, particularly in complex scenes, while increasing prompt length can enhance the model’s detection capabilities, it may also introduce more errors. A similar trend is observed in Figs. 8(b) and 8(d).In Fig. 8(c), when the prompt length is 4 (first row), the model successfully detects the correct object label (“Banana: 0.5610”) but also produces false labels, such as “Dining Table: 0.5085.” This indicates that, with shorter prompts, the model can correctly identify some objects while still suffering from false detections. However, when the prompt length increases to 8, 16, and 32 (second to fourth rows), the detection results exhibit notable errors, including high-confidence false labels like “Chair: 0.8581,” “TV monitor: 0.9042,” and “Dining Table: 0.7844.” These results suggest that in complex scenes containing numerous small objects, shorter prompt lengths help maintain the model’s object recognition capabilitie. In contrast, longer prompt lengths may incorporate additional features that lead to a significant increase in the false detection rate.
@@ IV-G Limitation and Future Work
While SCINet demonstrates state-of-the-art performance in PML tasks, it still presents certain limitations. Visual analysis reveals that increasing the length of learnable prompts can enhance detection capabilities; however, this often comes at the cost of a higher false detection rate, particularly in complex scenes with numerous small objects. This suggests a trade-off. Longer prompts may introduce additional, potentially noisy features, thereby undermining the model’s robustness and generalization ability, especially in scenarios characterized by occlusions or background clutter.
@@ V Conclusions
In this study, we introduce SCINet, a novel framework designed to tackle the challenges of PML tasks. At the core of SCINet is the use of a structured semantic prior from multi-modal representation learning to compensate for the lack of supervisory information by introducing a semantic co-occurrence insight network. SCINet combines text and visual features through a cross-modal fusion module, enhancing the model’s contextual understanding, feature extraction efficiency, and classification accuracy. Moreover, through the intrinsic semantic augmentation strategy, SCINet not only learns the features of the data but also captures its semantic meanings. This enables the model to excel in handling rare or ambiguous categories, improving its ability to generalize and manage the complexity and diversity of the real world.